### v2

# VidGuard-R1: AI-Generated Video Detection and Explanation via Reasoning MLLMs and RL

Abstract
With the rapid advancement of AI-generated videos, there is an urgent need for effective detection tools to mitigate societal risks such as misinformation and reputational harm. In addition to accurate classification, it is essential that detection models provide interpretable explanations to ensure transparency for regulators and end users. To address these challenges, we introduce VidGuard-R1, the first video authenticity detector that fine-tunes a multi-modal large language model (MLLM) using group relative policy optimization (GRPO). Our model delivers both highly accurate judgments and insightful reasoning. We curate a challenging dataset of 140k real and AI-generated videos produced by state-of-the-art generation models, carefully designing the generation process to maximize discrimination difficulty. We then fine-tune Qwen-VL using GRPO with two specialized reward models that target temporal artifacts and generation complexity. Extensive experiments demonstrate that VidGuard-R1 achieves state-of-the-art zero-shot performance on existing benchmarks, with additional training pushing accuracy above 95%. Case studies further show that VidGuard-R1 produces precise and interpretable rationales behind its predictions. The code is publicly available at https://VidGuard-R1.github.io.
Figure 1: Overall framework of VidGuard-R1. We present the first video authenticity detector based on multi-modal large language models (MLLMs), which generates a chain-of-thought reasoning process along with the final answer.
1 Introduction
In the past year, we have witnessed unprecedented progress in video generation models, with dramatic improvements in realism and quality. The release of powerful models such as Sora (Brooks et al., 2024), Wan (Wang et al., 2025a), and HunyuanVideo (Kong et al., 2024) has made AI-generated videos more accessible to the public, further blurring the line between synthetic videos and real ones. At the same time, these advancements have led to a series of social risks, including the spread of misinformation, violations of privacy rights, damage to personal reputations, and increased susceptibility to scams and fraud.
Motivated by its practical significance, several pioneering works have been developed to detect AI-generated videos. Early approaches primarily targeted DeepFake-style facial forgeries (Qian et al., 2020b; Tan et al., 2024; Gu et al., 2021), which often assumed single-subject, frontal-face scenarios under constrained settings. These assumptions diverge significantly from open-domain, multi-scene videos produced by modern generative models. More recent detectors leverage spatial-temporal consistency (Ma et al., 2024; Bai et al., 2024b; Liu et al., 2024); however, such methods are limited in capturing higher-level semantic or causal inconsistencies and can be easily bypassed by post-processing techniques. Other methods are trained on curated fake video detection datasets (Chen et al., 2024a; Ni et al., 2025; Kundu et al., 2025), but these benchmarks often lack coverage of newly emerging models and fail to reflect the full diversity of generative capabilities. A recent benchmark (Chen et al., 2024a) shows that even state-of-the-art detectors still struggle to reliably identify videos from advanced models like Sora. Furthermore, these detectors typically offer only binary decisions without accompanying explanations, which raises concerns for transparency, especially when detection outcomes affect content moderation or legal accountability. Users are also more likely to trust detection systems that provide interpretable reasoning.
Recent advances in multi-modal large language models (MLLMs) have significantly enhanced video understanding, enabling detailed explanations of model decisions (Bai et al., 2023; Zhang et al., 2024b). This makes them promising candidates for detecting and explaining AI-generated videos. However, directly applying existing MLLMs, including advanced models like GPT-4o, yields subpar performance on current benchmarks, underscoring the need for supervised fine-tuning (SFT). As an initial step, we applied SFT to the Qwen2.5-VL-7B model (Bai et al., 2025). While the model achieved strong overall performance, it remained limited in its ability to explain why a video is fake, revealing shortcomings in its reasoning capability.
To address this, we adopt reinforcement learning (RL), which has shown promise in enhancing LLM reasoning (Guo et al., 2025). Notably, Video-R1 (Feng et al., 2025) outperforms commercial models on video reasoning tasks. RL enables MLLMs to develop self-improving reasoning via outcome-based rewards. We hypothesize that RL fine-tuning can help models detect subtle temporal and generative artifacts. Key to this is designing effective reward models. Simple binary rewards (e.g., 1 for real, 0 for fake) are insufficient. Instead, we propose two strategies: (1) injecting temporal artifacts into both real and fake videos to encourage temporal reasoning, and (2) assigning higher rewards to videos generated with more diffusion steps, which are harder to detect. Incorporating these into a group relative policy optimization (GRPO) framework leads to over 86% accuracy on our dataset and 95% accuracy on two benchmarks.
  • We introduce VidGuard-R1, the first video authenticity detector that fine-tunes the MLLM using GRPO. The model leverages the pretrained knowledge of MLLMs for accurate classification and employs reinforcement learning for effective exploration. To further enhance performance, we design two specialized reward models that target temporal artifacts and generation complexity based on diffusion steps.
  • We construct a challenging dataset of 140k real/fake video pairs for AI-generated video detection. By employing state-of-the-art generation models and carefully controlling the process, we ensure that distinguishing real from fake is non-trivial.
  • VidGuard-R1 achieves state-of-the-art zero-shot performance on existing benchmarks, with accuracy exceeding 95%. Case studies further highlight its ability to produce accurate and interpretable explanations.
2 Related works
2.1 AI-generated video detection method
In recent years, research on AI-generated video detection has primarily focused on deepfake videos featuring synthetic faces (Pei et al., 2024), with methods based on spatial-temporal consistency, frequency artifacts, and data-driven approaches. However, these approaches often struggle to generalize beyond face-centric content to more diverse, real-world videos. Recently, general video detection methods have emerged: AIGDet (Bai et al., 2024a) captures spatial-temporal anomalies, DeCoF (Ma et al., 2024) exploits frame consistency, and diffusion-based representations have been used to track temporal dynamics (Liu et al., 2024). Other works identify key factors such as appearance, motion, and geometry for classifier training (Chang et al., 2024). MLLMs have also been explored for visual forgery detection: FakeShield (Xu et al., 2024) employs supervised fine-tuning (SFT) for image forgery detection and relies on GPT-4o for explanation generation, while SafeWatch (Chen et al., 2024b) applies both SFT and direct preference optimization (DPO) to train a video guardrail model with transparent reasoning. In contrast, our work is the first to fine-tune a multi-modal large language model (MLLM) using group relative policy optimization (GRPO) for AI-generated video detection, demonstrating strong generalization across a wide range of recent video generative models and benchmark datasets.
2.2 AI-generated video detection dataset
As this is a very recent area of research, only a few benchmarks have been proposed. The generated video dataset (GVD) (Bai et al., 2024a) (11k samples) and GenVideo (Chen et al., 2024a) (with millions of samples) consider settings where both training and test videos are generated by the same series of models. However, these benchmarks lack prompt/image–video pairs, semantic labels, or cross-source settings. GVF (2.8k samples) contains prompts/images–video pairs and semantic labels, but does not provide cross-source settings. GenVidBench (Ni et al., 2025) consists of 100k videos and incorporates cross-source settings, but the video generation models used are less advanced, such as CogVideo and SVD. To address these limitations, we construct a curated dataset of 140,000 real–fake video pairs generated with state-of-the-art video generation models (HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024)), standardizing video properties to mitigate superficial cues and encourage models to focus on intrinsic visual realism.
3 Methodology
Figure 2 illustrates the VidGuard-R1 framework, which consists of two stages. We first apply supervised fine-tuning (SFT) to the multimodal large language model (MLLM), followed by direct preference optimization (DPO) and group relative policy optimization (GRPO) based on the collected datasets. We further develop two GRPO variants by introducing temporal artifacts and leveraging videos generated with varying diffusion steps.
Figure 2: The overall training framework of VidGuard-R1, consisting of two stages: (1) supervised fine-tuning (SFT) for chain-of-thought (CoT) initialization, and (2) reinforcement learning-based fine-tuning to enable deeper reasoning.
3.1 Data collection
3.1.1 Data construction for video realism discrimination
High-quality training data is essential for video reasoning in MLLMs.. However, many existing benchmarks for real vs. generated video classification, such as GenVideo (Chen et al., 2024a) and GenVidBench  (Ni et al., 2025), exhibit uncontrolled discrepancies in basic metadata—e.g., real videos are often longer than 10 seconds while generated ones are typically under 4 seconds in GenVideo. Moreover, they reveal clear modality-level gaps in motion dynamics and content contrasts between real and generated videos. These differences introduce unintended shortcuts, enabling models to rely on superficial cues like duration or resolution rather than actual visual realism. As a result, VidGuard-R1 attains over 96% accuracy on both GenVideo and GenVidBench by exploiting such artifacts. To mitigate this reward hacking behavior, we construct a curated dataset with standardized video properties, encouraging models to focus on intrinsic visual content.
We collect real videos from the InternVid (Wang et al., 2023c) and ActivityNet (Caba Heilbron et al., 2015) datasets and generate their corresponding fake counterparts using HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024). We specifically choose these two models because they support conditioning on both the first-frame image and a text description—an essential requirement for generating videos that are contextually aligned with their real counterparts. To achieve such alignment, we provide the generation models with the first frame of each real video along with a textual caption describing its content. For ActivityNet, which lacks native captions, we extract concise descriptions using Qwen2.5-VL 72B. This pairing strategy mitigates content-based biases and forces the model to reason over subtle visual details.
3.1.2 Collecting chain of thought (CoT) annotation
Eliciting deliberate, step-by-step reasoning in MLLMs requires high-quality CoT supervision. To this end, we leverage Qwen-2.5-VL (72B) to extract salient visual cues from each video and guide the model toward a deeper understanding. Specifically, we query the model with critical factors known to distinguish real from generated content—motion consistency, lighting consistency, texture artifacts, and physical plausibility violations. These targeted prompts encourage detailed reasoning grounded in visual evidence.
However, current MLLMs lack the capacity to reliably distinguish real from fake videos on their own. To compensate, we provide ground-truth labels during prompt construction and instruct the model to generate CoT rationales conditioned on the given label. While these rationales do not reflect genuine discrimination ability, they capture rich contextual cues—such as object interactions, background details, and lighting inconsistencies—that are highly informative. These CoT annotations serve as useful clues for subsequent reinforcement learning fine-tuning. For prompt templates used in CoT generation, please refer to our supplementary materials.
3.2 Supervised and RL Fine-tuning
We begin with SFT, where the model is trained to mimic the ground-truth reasoning process. Given a video $x$ and its annotation $y$ from the collected dataset, the model parameters $\theta$ are optimized by minimizing the negative log-likelihood $\mathcal{L}_{\text{SFT}}(\theta)=-\sum_{t=1}^{T}\log p_{\theta}(y_{t}\mid y_{<t},x)$. To align the model outputs with human preferences, we apply DPO, which updates the model based on pairwise preference data without explicit reward modeling. Given a preferred response $y_{w}$ and a less-preferred response $y_{l}$ for the same video $x$, the DPO loss encourages the model to prefer $y_{w}$ over $y_{l}$ compared to a reference model $p_{\text{ref}}$:
$\mathcal{L}_{\text{DPO}}(\theta)=$ $-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\Bigg[\log\sigma\Big(\beta\log\frac{p_{\theta}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}-\beta\log\frac{p_{\theta}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\Big)\Bigg]$
where $\sigma(\cdot)$ is the sigmoid function and $\beta$ controls the preference strength. This method allows fine-tuning using preference comparisons without requiring scalar rewards.
Finally, we adopt GRPO from DeepSeek R1 (Guo et al., 2025), which generalizes RLHF to group-level comparisons. Given a query video $x$ and a group of generated outputs $\{o_{i}\}_{i=1}^{G}$, the model is trained to assign higher probabilities to outputs with higher rewards. The GRPO objective is: $\mathcal{L}_{\text{GRPO}}(\theta)=$ $-\mathbb{E}_{(x,o_{1:G})\sim D}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Bigg(\frac{p_{\theta}(o_{i}|x)}{p_{\text{ref}}(o_{i}|x)}A_{i},\text{clip}\left(\frac{p_{\theta}(o_{i}|x)}{p_{\text{ref}}(o_{i}|x)},1-\epsilon,1+\epsilon\right)A_{i}\Bigg)-\beta D_{\text{KL}}\left(p_{\theta}\,\|\,p_{\text{ref}}\right)\Bigg]$ where $\epsilon$ is a clipping threshold and $\beta$ regularizes the policy to stay close to the reference model. The advantage term $A_{i}$ normalizes the reward $r_{i}$ for output $o_{i}$ within the group, computed as $A_{i}=\tfrac{r_{i}-\mu_{x}}{\sigma_{x}}$, where $\mu_{x}$ and $\sigma_{x}$ are the mean and standard deviation of $\{r_{i}\}_{i=1}^{G}$. GRPO thus enables learning from relative ranking among multiple responses, capturing nuanced distinctions in quality across outputs.
3.3 VidGuard-R1
3.3.1 Overview
Figure 2 illustrates the training pipeline of VidGuard-R1. Following the data collection procedure, we construct two datasets of different scales: VidGuard-R1-CoT-30k and VidGuard-R1-RL-100k. We adopt Qwen2.5-VL-7B as the base MLLM and train it using our proposed fine-tuning framework.
The first stage is supervised fine-tuning initialization using the VidGuard-R1-CoT-30k dataset, which contains videos paired with chain-of-thought (CoT) annotations. This stage establishes foundational reasoning ability and equips the model with basic cross-modal alignment and visual understanding. The resulting model is referred to as VidGuard-R1 (CoT).
In the second stage, we apply two reinforcement learning methods—DPO and GRPO—to further refine the model on a larger and more diverse dataset, VidGuard-R1-RL-100k. DPO aligns the model with high-quality preference signals via pairwise comparisons, requiring the construction of preference pairs. Specifically, since our dataset includes pairwise real and fake videos, each sample is annotated with CoT rationales for both perspectives. For DPO training, we construct preference pairs by swapping these CoTs. For a real video, the CoT supporting its authenticity with the answer ”real” serves as the positive annotation, while the CoT from the paired fake video with the answer ”fake” is used as the negative annotation. In contrast, GRPO encourages consistent performance across grouped outputs by leveraging structural regularization. As it does not rely on preference annotations, video labels are directly used as reward signals. The resulting models are denoted as VidGuard-R1 (DPO) and VidGuard-R1 (GRPO).
We introduce two variants, GRPO-TA and GRPO-Q, to further enhance detection performance. These methods extend the original GRPO framework by adjusting reward values according to the difficulty of detecting fake videos. Detailed descriptions are provided in the following sections.
3.3.2 GRPO with temporal artifacts (GRPO-TA)
While standard GRPO performs well in video discrimination by leveraging local visual cues—such as pixel distortions and lighting inconsistencies—it often overlooks temporal inconsistencies, which are crucial for detecting generated videos. To address this limitation, we introduce GRPO with temporal artifacts (GRPO-TA), a variant that explicitly promotes temporal reasoning through a contrastive reward adjustment.
We apply two common temporal artifacts: (1) repeating a specific video segment and (2) reversing the frame sequence within a segment. These manipulations are applied probabilistically, with the manipulated region selected based on a Gaussian distribution over the video timeline.
Specifically, for each input query, we generate two sets of model outputs: $\{o_{i}\}_{i=1}^{G}$ for the original video, and $\{\tilde{o}_{i}\}_{i=1}^{G^{\prime}}$ for the corresponding manipulated video with temporal artifacts. These videos should be classified as fake videos. In GRPO-TA, we assign additional rewards when the model correctly classifies temporally manipulated videos as fake. Consider two numbers, $\alpha_{1}>\alpha_{2}$. Detecting temporal artifacts in videos manipulated from real content tends to be more challenging than identifying those derived from fake videos. This is because real videos typically exhibit coherent and natural motion, so temporal manipulations such as frame shuffling or repetition can be subtle and difficult to detect. In contrast, generated videos often contain artifacts like unstable motion or low temporal consistency, which make further manipulations more visually salient. To reflect this asymmetry in difficulty, we assign the model a higher reward $\alpha_{1}$ when the original video $o_{i}$ is real, and a moderate reward $\alpha_{2}$ when the original video is fake. This is defined as:
$w_{i}=\begin{cases}\alpha_{1},&\text{if }\tilde{o}_{i}=\textit{fake}\text{ and }y_{i}=\textit{real}\\ \alpha_{2},&\text{if }\tilde{o}_{i}=\textit{fake}\text{ and }y_{i}=\textit{fake}\\ 0,&\text{otherwise}\end{cases}$ (1)
where $y_{i}$ denotes the label of the $i$-th video. In the experiments, we set the hyperparameters to $\alpha_{1}=0.5$ and $\alpha_{2}=0.3$. This additional reward, $w_{i}$, is designed to be applied conditionally. Specifically, for a given sample, we only add $w_{i}$ to the original GRPO reward if two conditions are met: the model’s prediction on the original video ($O_{i}$) must be correct, and the overall accuracy on the group of manipulated videos ($\tilde{p}$) must exceed a predefined threshold $\mu$. This ensures that we only reward the model for temporal reasoning when it already has a solid baseline performance. The final reward of GRPO-TA is given by
$r_{i}^{\text{GRPO-TA}}=\begin{cases}r_{i}^{\text{GRPO}}+w_{i},&\text{if }o_{i}\text{ is correct and }\tilde{p}>\mu\\ r_{i}^{\text{GRPO}},&\text{otherwise}\end{cases}$ (2)
where $r_{i}$ denotes the original GRPO reward, set to 1 if the model prediction on the original video is correct, and 0 otherwise. The additional reward $w_{i}$ is applied only when both the original prediction is correct and the group of responses for the temporally manipulated videos achieves higher accuracy. In the experiments, we set $\mu=0.8$.
3.3.3 GRPO with quality evolutionary videos (GRPO-Q)
Our motivation is to extend the model’s capability to detect videos based on quality. Given the subjective nature of quality assessment, we avoid relying on large-scale human annotations. Instead, we leverage diffusion-based video generation by systematically varying the number of reverse diffusion steps to produce videos with distinct quality levels.
As in GRPO-TA, $o_{i}\in\mathcal{Y}$ and $y_{i}\in\mathcal{Y}$ denote the model output and ground-truth label, with $\mathcal{Y}=\{\text{real}\}\cup\{\text{fake-}s\}$, where $s$ is the diffusion step. A reward is given for an exact match, and no reward is assigned if the real/fake classification is incorrect. In GRPO-Q, if the model correctly classifies a fake video but selects an incorrect diffusion step, we assign a partial reward based on the distance between the predicted and ground-truth diffusion steps. The GRPO-Q reward is defined as follows:
[…]
The first scenario occurs when the model fails to correctly classify the video as real or fake. The second scenario, where $\delta=1$, represents an exact match in prediction, including the diffusion progression. In the third case, the function $g(\cdot,\cdot)$ maps the step distance to a scalar reward, enabling fine-grained credit assignment based on the similarity in quality. Specifically, we define a progress value $s()$ in the range $[0,1]$ to indicate the fraction of diffusion steps used, where 0 denotes zero steps, and $1$ denotes full completion of the steps. The ground-truth value is $s(y_{i})$, and the model will estimate a progress value. We define the reward function as $g(o_{i},y_{i})=\delta\cdot(1-|s(o_{i})-s(y_{i})|)$.
This reward formulation enables the model to move beyond binary discrimination and perform fine-grained analysis of video quality. By learning to associate subtle differences in generation steps with quality variations, the model develops a deeper understanding of the diffusion process and its impact on perceptual realism. As a result, it can not only detect whether a video is fake, but also infer but also estimate the degree of quality degradation in generated videos. This facilitates more interpretable and controllable evaluation of generated content quality.
4 Experiments
4.1 Implementation details
4.1.1 Dataset
Our dataset contains 140k videos, balanced between 70k real and 70k generated samples, organized into contextual pairs. The real set comprises 55k videos from InternVid and 15k from ActivityNet, while the generated set includes 50k samples synthesized by HunyuanVideo-I2V (Kong et al., 2024) and 20k by CogVideoX-5B (Yang et al., 2024). We allocate 130k samples for training and 10k for testing, with the latter evenly split between real and generated videos. Within the training data, 30k samples are reserved for chain-of-thought (CoT) learning, denoted as VidGuard-R1-CoT-30k, and the remaining 100k are used for reinforcement learning fine-tuning, denoted as VidGuard-R1-RL-100k.
[…]
For GRPO-Q fine-tuning, we augment the training set with intermediate generations sampled from diffusion steps 10 to 50. These are labeled with approximate quality levels (20%, 40%, 60%, 80%, and 95%). Specifically, we use 12k real videos, each paired with five generated variants at different diffusion steps, resulting in 72k samples per generation model.
4.1.2 Evaluation protocol
We evaluate three datasets—ours, GenVidBench (Ni et al., 2025), and GenVideo (Chen et al., 2024a)—using the metrics and baselines defined by their respective benchmarks. For ours and GenVidBench, we report mean Top-1 accuracy, the average correctness over all predictions. For GenVideo, we follow the original protocol and report recall and F1 score. All evaluations adhere to the official settings of each benchmark to ensure fair comparison.
4.1.3 Training Setup
We employ Qwen2.5-VL-7B as the base MLLM and conduct all experiments on four NVIDIA A100 GPUs (80GB). Each video is represented by up to 16 frames, where each frame is resized to a $28\times 28$ spatial resolution and mapped to 128 feature channels for encoder input during both training and inference. For GenVideo and GenVidBench, we follow their official evaluation protocols and adopt 8-frame inputs. In GRPO training, we sample 8 responses per input; for GRPO-TA, we additionally sample 4 responses from temporally manipulated variants of the input to enhance robustness against temporal artifacts. Training proceeds in two stages: first, the base model is fine-tuned for one epoch on the CoT dataset, yielding the SFT-CoT MLLM; second, we initialize VidGuard-R1 with SFT-CoT and perform reinforcement learning for approximately 2,000 steps.
Table 1: Comparison of models on our dataset, reported as mean Top-1 accuracy (%). TF denotes transformer.
Method Type CogVideoX HunyuanVideo
SlowFast CNN 77.87 77.03
I3D CNN 64.78 62.13
TRN CNN 68.73 69.87
UniFormer V2 TF 73.95 71.92
TimeSformer TF 78.53 74.55
VideoSwin TF 76.81 79.71
MViT V2 TF 58.38 53.91
Qwen2.5-VL-7B MLLM 50.95 52.83
GPT-4.1 mini MLLM 54.95 55.31
VidGuard-R1 (CoT) MLLM 66.18 63.19
VidGuard-R1 (DPO) MLLM 79.13 80.88
VidGuard-R1 (GRPO) MLLM 81.30 81.90
VidGuard-R1 (GRPO-TA) MLLM 82.17 83.72
VidGuard-R1 (GRPO-Q) MLLM 84.32 86.17
Table 2: Extended GenVidBench results with VidGuard-R1 and additional MLLMs, reported as mean Top-1 accuracy (%). TF denotes transformer.
Method Type MuseV SVD CogVideo Mora HD-VG Mean
SlowFast CNN 12.25 12.68 38.34 45.93 93.63 41.66
I3D CNN 8.15 8.29 60.11 59.24 93.99 49.23
TRN CNN 38.92 26.64 91.34 93.98 93.97 71.26
UniFormer V2 TF 20.05 14.81 45.21 99.21 96.89 57.55
TimeSformer TF 73.14 20.17 74.80 39.40 92.32 64.28
VideoSwin TF 62.29 8.01 91.82 45.83 99.29 67.27
MViT V2 TF 76.34 98.29 47.50 96.62 97.58 79.90
Qwen2.5-VL-7B MLLM 25.86 27.06 68.51 43.26 71.15 47.30
GPT-4.1 mini MLLM 26.07 33.78 94.07 57.19 87.64 59.62
VidGuard-R1 (CoT) MLLM 36.52 16.02 99.35 76.94 99.94 66.09
VidGuard-R1 (GRPO,
GenVideo-pretrained,
Zero-shot)
MLLM 97.24 96.59 99.88 99.93 88.14 96.37
VidGuard-R1 (GRPO) MLLM 97.38 94.98 99.90 99.99 95.46 97.53
4.2 Main results
4.2.1 Our dataset
We evaluate VidGuard-R1 on our dataset with several methods, including CNN-based models (SlowFast (Feichtenhofer et al., 2019), I3D (Carreira & Zisserman, 2017), TRN (Zhou et al., 2018)), Transformer-based models (UniFormer V2 Li et al. (2022a), TimeSformer (Bertasius et al., 2021), VideoSwin (Liu et al., 2022), MViT V2 (Li et al., 2022b)), and MLLM-based models (Qwen2.5-VL (Bai et al., 2025) and GPT-4.1 mini (OpenAI, 2025)). For CNN and Transformer models, we use the default training settings provided by the MMAction2 framework (Contributors, 2020).
As shown in Table 2, CNN- and Transformer-based models achieved 53–79% accuracy, with SlowFast and TimeSformer among the top performers. In contrast, Qwen2.5-VL-7B and GPT-4.1 mini exhibited near-random performance, highlighting their limited capability in distinguishing fake videos. VidGuard-R1 (CoT), trained via supervised fine-tuning (SFT) on Qwen2.5-VL-7B, substantially improved accuracy from around 51% to over 66%, yet remained less competitive compared to advanced SOTA methods. This result aligns with the intended role of the SFT stage—as a cold start phase to guide the model toward structured think + answer responses, emphasizing the extraction of salient visual cues.
In the subsequent RL stage, both DPO and GRPO further improved performance by roughly 2% over the best baseline. Our proposed methods—GRPO-TA and GRPO-Q—achieved additional gains of approximately 2% and 5% over GRPO, respectively, demonstrating the effectiveness of temporal artifact supervision and quality-aware reward modeling in enhancing detection accuracy.
4.2.2 GenVidBench benchmark
The GenVidBench dataset comprises approximately 87k training samples and 82k testing samples, with fake videos generated by models such as MuseV (Xia et al., 2024), SVD (Blattmann et al., 2023), CogVideo (Hong et al., 2022), and Mora (Yuan et al., 2024), and real videos sourced from HD-VG (Wang et al., 2023b). We conduct training and evaluation under the cross-source and cross-generator settings as proposed in their benchmark. In addition to the models originally reported in GenVidBench, we evaluate VidGuard-R1 using the same model families as in our dataset experiments—CNN-based, Transformer-based, and MLLM-based models—including two MLLMs: Qwen2.5-VL and GPT-4.1 mini. VidGuard-R1 (GRPO, GenVideo-pretrained, Zero-shot) denotes the zero-shot model pretrained on GenVideo and evaluated on GenVidBench. As shown in Table 2, both the zero-shot model and two fine-tuned variants achieve over 15% higher accuracy compared to prior SOTA methods. Notably, the zero-shot model demonstrates strong generalization, highlighting the effectiveness of pretraining on diverse generative content. Complete detection model results are provided in Appendix B.
Table 3: Extended GenVideo results with VidGuard-R1 and additional MLLMs, evaluated by F1 score and recall (R)
Method Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Mean
level Studio Valley Scope Scrape
NPR (Tan et al., 2024) Image R 0.91 0.99 0.99 0.24 0.89 0.57 0.97 0.99 0.94 0.87 0.84
F1 0.27 0.84 0.91 0.30 0.86 0.59 0.81 0.91 0.81 0.81 0.71
VideoMAE (Tong et al., 2022) Video R 0.67 0.96 0.98 0.96 0.77 0.80 0.97 0.96 0.96 0.68 0.87
F1 0.62 0.95 0.98 0.96 0.86 0.87 0.96 0.97 0.96 0.79 0.89
MINTIME-CLIP (Coccomini et al., 2024) Video R 0.89 1.00 0.98 0.26 0.96 0.98 0.99 1.00 0.84 0.82 0.87
F1 0.49 0.93 0.96 0.37 0.94 0.92 0.92 0.96 0.84 0.85 0.82
FTCN-CLIP (Zheng et al., 2021) Video R 0.87 1.00 0.98 0.17 0.97 0.91 1.00 1.00 0.85 0.82 0.86
F1 0.78 0.98 0.98 0.29 0.98 0.94 0.98 0.99 0.90 0.89 0.87
DeMamba-XCLIP (Chen et al., 2024a) Video R 0.98 1.00 0.99 0.65 0.94 0.98 1.00 1.00 0.92 0.89 0.93
F1 0.64 0.96 0.97 0.75 0.95 0.95 0.95 0.97 0.92 0.91 0.90
Qwen2.5-VL-7B (Bai et al., 2025) MLLM R 0.58 0.56 0.54 0.33 0.43 0.38 0.81 0.63 0.51 0.70 0.54
F1 0.74 0.72 0.70 0.49 0.60 0.55 0.90 0.77 0.68 0.82 0.70
GPT-4.1 mini (OpenAI, 2025) MLLM R 0.43 0.67 0.56 0.54 0.63 0.56 0.92 0.67 0.69 0.69 0.65
F1 0.60 0.80 0.72 0.70 0.77 0.72 0.96 0.80 0.82 0.82 0.72
VidGuard-R1 (CoT) MLLM R 0.92 0.89 0.91 0.90 0.98 0.79 0.99 0.85 0.89 0.87 0.90
F1 0.90 0.91 0.95 0.89 0.99 0.81 0.95 0.89 0.85 0.88 0.90
VidGuard-R1 (GRPO,
GenVidBench-pretrained, Zero-shot)
MLLM R 0.95 0.98 0.90 0.89 0.97 0.85 0.99 0.93 0.81 0.87 0.92
F1 0.93 0.93 0.96 0.91 0.99 0.82 0.95 0.89 0.85 0.88 0.91
VidGuard-R1 (GRPO) MLLM R 0.95 1.00 0.98 0.94 0.98 0.95 0.97 0.99 0.94 0.91 0.96
F1 0.97 0.99 0.99 0.91 0.99 0.89 0.99 0.99 0.95 0.90 0.96
4.2.3 GenVideo benchmark
The GenVideo dataset comprises approximately 2.2M training samples and 20k testing samples, with generated videos sourced from a diverse set of models, including Sora (OpenAI, 2024), MorphStudio (mor, 2025), Gen2 (Esser et al., 2023b), HotShot (hot, 2025), Lavie (Wang et al., 2025b), Show-1 (Zhang et al., 2024a), MoonValley (moo, 2025), Crafter (Chen et al., 2023), ModelScope (Wang et al., 2023a), and WildScrape (Wei et al., 2024). Following the official evaluation protocol, we benchmark two MLLM baselines and three variants of VidGuard-R1. Among these, VidGuard-R1 (GRPO) consistently outperforms almost all prior detection methods across videos generated by the various models. As shown in Table 3, it achieves an F1 score improvement of 0.06 compared to DeMamba-XCLIP. Complete detection model results are provided in Appendix B.
4.2.4 Performance gap between our dataset and benchmarks
While VidGuard-R1 achieves approximately 85% accuracy on our dataset, it obtains significantly higher accuracy—exceeding 95%—on the two benchmark datasets. This discrepancy arises from two key differences. First, the benchmarks exhibit clear discrepancies in video metadata—such as resolution, duration, and frame rate—between real and fake videos, which models can exploit as superficial cues. In contrast, we standardize all videos in our dataset by matching resolution, FPS, and format, thereby forcing models to rely on actual visual content. Second, our dataset ensures strong contextual alignment by conditioning generation on the first frame and the corresponding caption of a real video, resulting in more realistic and semantically consistent outputs. In comparison, benchmark datasets often generate fake videos from unrelated prompts and synthetic images, leading to artifacts that make detection easier.
4.2.5 Ablation study
Table 4: LLM-as-a-Judge explanation scores on our dataset
Method
Expl. Score
Expl. Score
Qwen2.5-VL-7B 5.8 5.6
GPT-4.1 mini 5.8 5.9
VidGuard-R1 (CoT) 6.8 6.9
VidGuard-R1 (DPO) 7.2 8.1
VidGuard-R1 (GRPO) 8.1 8.0
VidGuard-R1 (GRPO-TA) 8.1 8.4
VidGuard-R1 (GRPO-Q) 8.3 8.5
Explanation quality and accuracy comparison
Table 4 presents results on the HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024) datasets. We report explanation quality scores, which are rated on a 1–10 scale (with 10 indicating excellent quality and full alie code is publicly availagnment) by GPT-4.1 mini using the LLM-as-a-Judge prompt described in Appendix C. Compared to baseline models such as Qwen2.5-VL-7B and GPT-4.1 mini, our VidGuard-R1 GRPO variants achieve consistent improvements in both classification accuracy and explanation quality.
Table 5: Accuracy (%) for GRPO-TA under different reward function parameters $\alpha_{1}$ and $\alpha_{2}$
$\alpha_{1}$ $\alpha_{2}$ Accuracy (%)
0.3 0.1 81.31
0.3 0.3 82.59
0.5 0.3 83.57
0.5 0.5 83.12
0.7 0.5 82.53
GRPO-TA reward ablation
Table 5 reports an ablation study of GRPO-TA on our dataset by varying the weight parameters $\alpha_{1}$ and $\alpha_{2}$, which control the relative importance of different temporal artifact types. The highest classification accuracy of 83.57% is achieved with $\alpha_{1}=0.5$ and $\alpha_{2}=0.3$, while the threshold $\mu$ is fixed at 0.8 across all experiments.
Table 6: Accuracy (%) for GRPO-Q with varying number of intermediate diffusion steps
# of steps (step numbers) Accuracy (%)
1 (50) 81.63
3 (10, 30, 50) 83.21
5 (10, 20, 30, 40, 50) 85.80
GRPO-Q reward ablation
Table 6 presents an ablation study on GRPO-Q conducted on our dataset by varying the number of intermediate diffusion steps included per real video during fine-tuning. Using more steps provides richer supervision of video quality progression, improving detection accuracy. The best accuracy of 84.05% is obtained with five steps, which is the setting used in our main experiments.
4.3 Case studies on explanations
Figure 3: VidGuard-R1 (GRPO): reasoning about an origami folding sequence
Figure 4: VidGuard-R1 (GRPO-Q): temporal inconsistency observed as unnatural movements
Figures 4 and 4 illustrate cases where VidGuard-R1 correctly identifies videos as generated. The model performs multi-faceted reasoning across motion, lighting, texture, and physical plausibility before arriving at a final decision. Notably, it does not rely on a single cue, but instead accumulates evidence across frames, resembling how humans distinguish fake videos. In each figure, pink highlights denote cues suggesting realism, red indicates artifacts indicative of generation, yellow marks intermediate reasoning steps, and underlines represent several key factors.
For instance, in Figure 4, the smooth hand motion initially suggests realism; however, once the origami folds in a physically implausible manner, the model revises its judgment. In Figure 4, although the lighting and shadows are consistent—typically a cue for authenticity—the model recognizes that this is insufficient in a largely static scene with only a stationary machine and a human hand. In particular, even in its final prediction, the model reflects on earlier realistic cues and acknowledges that the video could still be real, underscoring its nuanced, human-like reasoning in assessing video quality. Additional case studies are provided in the Appendix D.
5 Conclusion
We propose VidGuard-R1, an MLLM-based discriminator that not only detects AI-generated videos with high accuracy but also provides interpretable reasoning. By leveraging reinforcement learning with reward models targeting temporal artifacts and generation quality, VidGuard-R1 achieves 85% accuracy on our dataset, 97% on GenVidBench, and 96% on GenVideo, substantially surpassing prior state-of-the-art methods. We expect this work to advance MLLM-based video analysis and foster future research on strengthening MLLMs’ reasoning.
5.1 Limitations
Our dataset currently includes fake videos generated exclusively with the HunyuanVideo and CogVideoX. Although it is constructed in a pairwise fashion to ensure contextual similarity between real and fake videos, its generalizability remains limited. Expanding the dataset to incorporate fake videos from a broader range of generative models would result in a more diverse and robust training set, thereby improving its applicability to real-world scenarios.
Appendix A Additional Setup
To further guide the model during RL training, we incorporate a length-based reward strategy. We promote informative yet concise reasoning by rewarding outputs that are neither too brief nor excessively long. Specifically, if the model predicts the correct answer and the length of the response falls within the range $l_{\min},l_{\max}]$, an additional reward $\omega$ is assigned. Let $l_{i}$ be the length of the model’s response for the $i$-th video. The reward is defined as:
$r_{i}^{total}=\begin{cases}r_{i}+\omega,&\text{if }o_{i}\text{ is correct and }l_{\min}\leq l_{i}\leq l_{\max}\\ r_{i},&\text{otherwise}\end{cases}$ (4)
where we set $\omega=0.1$, $l_{\min}=320$, and $l_{\max}=512$.
Appendix B Comprehensive Benchmark Evaluation
In this section, we provide extended benchmark results for VidGuard-R1 alongside additional MLLMs. Table 7 presents mean Top-1 accuracy on GenVidBench across multiple video datasets, including CNN and Transformer baselines as well as selected MLLM variants. Table 8 reports comprehensive F1 and recall scores on the GenVideo dataset, including all models provided in the official benchmark alongside our MLLM variants. These extended tables offer a complete comparison of performance across all evaluated models.
Table 7: Extended GenVidBench results with VidGuard-R1 and additional MLLMs, reported as mean Top-1 accuracy (%). TF denotes transformer.
Method Type MuseV SVD CogVideo Mora HD-VG Mean
SlowFast (Feichtenhofer et al., 2019) CNN 12.25 12.68 38.34 45.93 93.63 41.66
F3Net (Qian et al., 2020a) CNN 37.43 37.27 36.46 39.59 52.76 42.52
I3D (Carreira & Zisserman, 2017) CNN 8.15 8.29 60.11 59.24 93.99 49.23
CFV2 (Nguyen et al., 2019) CNN 86.26 86.53 10.10 16.90 88.40 60.53
TPN (Yang et al., 2020) CNN 37.86 8.79 68.25 90.04 97.34 61.52
TIN (Shao et al., 2020) CNN 33.78 21.47 81.59 79.44 97.88 63.97
TRN (Zhou et al., 2018) CNN 38.92 26.64 91.34 93.98 93.97 71.26
TSM (Lin et al., 2019) CNN 70.37 54.70 78.46 70.37 96.76 76.40
X3D (Feichtenhofer, 2020) CNN 92.39 37.27 65.72 49.60 97.51 77.09
UniFormer V2 (Li et al., 2022a) TF 20.05 14.81 45.21 99.21 96.89 57.55
TimeSformer (Bertasius et al., 2021) TF 73.14 20.17 74.80 39.40 92.32 64.28
VideoSwin (Liu et al., 2022) TF 62.29 8.01 91.82 45.83 99.29 67.27
MViT V2 (Li et al., 2022b) TF 76.34 98.29 47.50 96.62 97.58 79.90
Qwen2.5-VL-7B (Bai et al., 2025) MLLM 25.86 27.06 68.51 43.26 71.15 47.30
GPT-4.1 mini (OpenAI, 2025) MLLM 26.07 33.78 94.07 57.19 87.64 59.62
VidGuard-R1 (CoT) MLLM 36.52 16.02 99.35 76.94 99.94 66.09
VidGuard-R1 (GRPO,
GenVideo-pretrained, Zero-shot)
MLLM 97.24 96.59 99.88 99.93 88.14 96.37
VidGuard-R1 (GRPO) MLLM 97.38 94.98 99.90 99.99 95.46 97.53
Table 8: Extended GenVideo results with VidGuard-R1 and additional MLLMs, evaluated by F1 and recall scores
Model Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Mean
level Studio Valley Scope Scrape
F3Net (Qian et al., 2020a) Image R 0.83 0.99 0.98 0.77 0.57 0.36 0.99 0.99 0.89 0.76 0.81
F1 0.50 0.94 0.96 0.81 0.69 0.49 0.93 0.96 0.88 0.82 0.80
NPR (Tan et al., 2024) Image R 0.91 0.99 0.99 0.24 0.89 0.57 0.97 0.99 0.94 0.87 0.84
F1 0.27 0.84 0.91 0.30 0.86 0.59 0.81 0.91 0.81 0.81 0.71
STIL (Gu et al., 2021) Video R 0.78 0.98 0.98 0.76 0.61 0.53 0.99 0.97 0.94 0.65 0.82
F1 0.38 0.90 0.94 0.78 0.72 0.62 0.90 0.94 0.88 0.72 0.78
VideoMAE (Tong et al., 2022) Video R 0.67 0.96 0.98 0.96 0.77 0.80 0.97 0.96 0.96 0.68 0.87
F1 0.62 0.95 0.98 0.96 0.86 0.87 0.96 0.97 0.96 0.79 0.89
MINTIME-CLIP (Coccomini et al., 2024) Video R 0.89 1.00 0.98 0.26 0.96 0.98 0.99 1.00 0.84 0.82 0.87
F1 0.49 0.93 0.96 0.37 0.94 0.92 0.92 0.96 0.84 0.85 0.82
FTCN-CLIP (Zheng et al., 2021) Video R 0.87 1.00 0.98 0.17 0.97 0.91 1.00 1.00 0.85 0.82 0.86
F1 0.78 0.98 0.98 0.29 0.98 0.94 0.98 0.99 0.90 0.89 0.87
TALL (Xu et al., 2023) Video R 0.91 0.98 0.97 0.83 0.76 0.79 0.99 0.98 0.94 0.66 0.88
F1 0.26 0.82 0.89 0.74 0.77 0.72 0.81 0.90 0.80 0.67 0.74
CLIP (Radford et al., 2021) Image R 0.94 0.99 0.91 0.77 0.88 0.86 0.99 0.99 0.84 0.84 0.90
F1 0.28 0.84 0.86 0.72 0.85 0.76 0.82 0.91 0.76 0.79 0.76
DeMamba-CLIP (Chen et al., 2024a) Video R 0.95 1.00 0.98 0.69 0.92 0.93 1.00 1.00 0.83 0.82 0.91
F1 0.64 0.96 0.97 0.78 0.94 0.92 0.95 0.98 0.87 0.87 0.89
XCLIP (Ni et al., 2022) Video R 0.82 0.99 0.93 0.61 0.79 0.69 0.97 0.99 0.77 0.83 0.84
F1 0.31 0.88 0.90 0.65 0.82 0.70 0.86 0.93 0.75 0.82 0.76
DeMamba-XCLIP (Chen et al., 2024a) Video R 0.98 1.00 0.99 0.65 0.94 0.98 1.00 1.00 0.92 0.89 0.93
F1 0.64 0.96 0.97 0.75 0.95 0.95 0.95 0.97 0.92 0.91 0.90
Qwen2.5-VL-7B (Bai et al., 2025) MLLM R 0.58 0.56 0.54 0.33 0.43 0.38 0.81 0.63 0.51 0.70 0.54
F1 0.74 0.72 0.70 0.49 0.60 0.55 0.90 0.77 0.68 0.82 0.70
GPT-4.1 mini (OpenAI, 2025) MLLM R 0.43 0.67 0.56 0.54 0.63 0.56 0.92 0.67 0.69 0.69 0.65
F1 0.60 0.80 0.72 0.70 0.77 0.72 0.96 0.80 0.82 0.82 0.72
VidGuard-R1 (CoT) MLLM R 0.92 0.89 0.91 0.90 0.98 0.79 0.99 0.85 0.89 0.87 0.90
F1 0.90 0.91 0.95 0.89 0.99 0.81 0.95 0.89 0.85 0.88 0.90
VidGuard-R1 (GRPO,
GenVidBench-pretrained, Zero-shot)
MLLM R 0.95 0.98 0.90 0.89 0.97 0.85 0.99 0.93 0.81 0.87 0.92
F1 0.93 0.93 0.96 0.91 0.99 0.82 0.95 0.89 0.85 0.88 0.91
VidGuard-R1 (GRPO) MLLM R 0.95 1.00 0.98 0.94 0.98 0.95 0.97 0.99 0.94 0.91 0.96
F1 0.97 0.99 0.99 0.91 0.99 0.89 0.99 0.99 0.95 0.90 0.96
Appendix C Prompt
Figure 5 shows the base prompt used for the real-vs-fake classification task. Annotators are instructed to assess whether a video is real or AI-generated by analyzing key visual and physical properties.
Figures 6 and 7 provide category-specific rationale collection prompts. In particular, Figure 6 presents the prompt for identifying visual cues of realism in real videos, while Figure 7 focuses on spotting artifacts in AI-generated videos. Both prompts guide annotators to evaluate videos across four diagnostic categories: motion consistency, lighting consistency, texture artifacts, and physics violations.
Figure 8 illustrates the LLM-as-a-Judge prompt used to evaluate rationale quality. In this setting, GPT-4.1 mini rates the quality of model-generated explanations on a 1–10 scale, where a score of 10 corresponds to excellent quality and full alignment with the ground truth rationale.
Figure 5: Prompt for identifying realism cues in real videos across four categories
Figure 6: Prompt for identifying realism cues in real videos across four categories
Figure 7: Prompt for identifying artifacts in AI-generated videos across four categories
Figure 8: LLM-as-a-Judge Prompt for Rationale Quality Evaluation
Appendix D Case studies on explanations
D.1 GenVidBench
Figures 912 present inference examples for videos synthesized by four distinct AI models included in the GenVidBench testing dataset: MuseV (Xia et al., 2024), SVD (Blattmann et al., 2023), CogVideo (Hong et al., 2022), and Mora (Yuan et al., 2024).
D.2 GenVideo
Figures 1322 show inference examples for videos generated by ten different AI models included in the GenVideo testing dataset: Sora (OpenAI, 2024), Morph Studio (mor, 2025), Gen2 (Esser et al., 2023a), HotShot (hot, 2025), Lavie (Wang et al., 2025b), Show-1 (Zhang et al., 2024a), Moonvalley (moo, 2025), Crafter (Chen et al., 2023), ModelScope (Wang et al., 2023a), and DreamVideo (Wei et al., 2024).
Figure 9: VidGuard-R1 (GRPO) inference on a MuseV-generated video
Figure 10: VidGuard-R1 (GRPO) inference on a SVD-generated video
Figure 11: VidGuard-R1 (GRPO) inference on a CogVideo-generated video
Figure 12: VidGuard-R1 (GRPO) inference on a Mora-generated video
Figure 13: VidGuard-R1 (GRPO) inference on a Sora-generated video
Figure 14: VidGuard-R1 (GRPO) inference on a Morph Studio-generated video
Figure 15: VidGuard-R1 (GRPO) inference on a Gen2-generated video
Figure 16: VidGuard-R1 (GRPO) inference on a HotShot-generated video
Figure 17: VidGuard-R1 (GRPO) inference on a Lavie-generated video
Figure 18: VidGuard-R1 (GRPO) inference on a Show-1-generated video
Figure 19: VidGuard-R1 (GRPO) inference on a Moonvalley-generated video
Figure 20: VidGuard-R1 (GRPO) inference on a Crafter-generated video
Figure 21: VidGuard-R1 (GRPO) inference on a ModelScope-generated video
Figure 22: VidGuard-R1 (GRPO) inference on a DreamVideo-generated video